Project

General

Profile

Actions

RunPlan20150714

Preparation for running and training shifters

Daily schedule

July 14
estimated time 8:00-12:00 14:00-18:00
MOM R. Bayes R. Bayes
Lead shifter Francois Domizia
Shifter 2 Domizia Galena
Shifter 3 Roumen
Shifter 4 Vittorio
BLOC Joe
SOC
Experimenter
Aim Move EMR, KL, TOF2 Training
Beam? No Yes
Additional Personnel Craig MacW
Henry

Beamline Configurations

We expect to be running with 2V beamloss and beam bump active. Beam line training will be conducted.

Run Time Preparedness.

  • Downstream PID Detectors (EMR/KL/TOF)
    • in final positions
    • HV on
  • Upstream PID Detectors (Ckov/TOFs)
    • in final positions -- Done (July 4)
  • Trackers
    • Helium on
      • downstream tracker -- Done (July 3)
      • Upstream tracker still has a leak
    • Repair dead channels.
  • Run Control
    • CDB interface
      • Need to improve stability of run control - remove need for IOC restart
    • CDB beam line control
      • Starting beam line current from zero doesn't work due to polarimeter -- Fixed (July 12)
  • Beam line control
    • "Set all to 0" button has not operated in a predictable manner -- Corrected. (July 12)
  • Beam line
    • Decay Solenoid on standby -- Done (July 13)
    • Correcting limits on the alarm handler.

Expected Work in the Hall

  • Movement of EMR/KL/TOF today
    • external company to provide the lift
    • will need equipment doors opened
    • Francois and Domizia should be on hand to reconnect the EMR and KL
  • Pipe welding
    • should have taken two days.

Phone details for 6pm daily meeting

Phone access : ReadyTalk number

US 866-740-1260

UK 0800-496-0576

Swiss 0800-70-53-51

Italy 800182592

Access Code: 2155402

Updated by Bayes, Ryan about 7 years ago ยท 5 revisions